OpenStudy (anonymous):

find the square roots of this polynomial: (4x^2)/(y^2)-(36x)/y+81

OpenStudy (anonymous):

\[\frac{ 4x^2 }{ x^2 }-\frac{ 36x}{ y }+81\]

OpenStudy (anonymous):

(4x²)/(y²) - (36x)/y + 81 = 0 0 = 4(x/y)² - 36(x/y) + 81 0 = [ 2(x/y) - 9 ] [ 2(x/y) - 9 ] ......... (x/y) = 9/2

OpenStudy (anonymous):

uinderstand?

OpenStudy (anonymous):

wow, i never would have gotten that answer

OpenStudy (anonymous):

not at all

OpenStudy (anonymous):

but i have it copied so i can look at it later,,,,

OpenStudy (anonymous):

k

OpenStudy (anonymous):

ty!

OpenStudy (anonymous):

np